of all at least one of t (a) What is the probability that a randomly selected adult regularly consumes both coffee and soda? (b) What is the probability that a randomly selected adult doesnt regularly consume at least one of these two products? Need Help? ReTalk to Titer
Answer #1

a)P(regularly consume both) =0.65+0.60-0.75=0.50

b)P(does not consume at least one of the products)=1-P(regularly consume both) =0.5